So i have used all day trying to figure out how I can determine the element out of a given function when asked "Which element has the following series of ionization energies I1, I2, I3 and I4?". Can you please explain briefly? With examples

This is a typical problem I will get. I only want explanation and steps for how to solve such problems. I only want explanation of how to solve problems like that, I dont have an example.
Answer #1

Ionization energy is the energy required when a atom in its complete gaseous form loses one electron from its most external orbital or valence shell. So for a element to have I1, I2, I3, I4 it needs to have at least 4 such electrons in its valence shell that it can lose.

Usually elements like Si, Ge, Sn who has ns2np2 electronic structure in its valence shell can lose electrons one by one to give I1, I2, I3 and I4 till they can get a stable octate in their most outer shell (valence shell).

For example:

Si ([Ne]3s23p2) - e = Si+  ([Ne]3s23p1) I1

Si ([Ne]3s23p1) - e = Si2+  ([Ne]3s2) I2

Si2+ ([Ne]3s2) - e = Si3+  ([Ne]3s1)​​​​​​​ I3

Si3+ ([Ne]3s1) - e = Si4+  ([Ne])​​​​​​​ I4

Similarly some of the transition metals can as well undergo such ionizations having (n-1)dxns2 electronic configurations where x is generally from 2 to 9. For an example Ti ([Xe]3d24s2) shows such type of ionization.